What is the change in chemical potential of glucose when its concentration in water at 20.0C is changed from 0.10 mol dm^-3 to 1.00 mol dm^-3?
1
Expert's answer
2017-03-09T09:43:06-0500
Δμ=RTLnC glucose1 / C glucose2 =8.314*273*Ln1/0.1=5226 J/mol